OpenStudy (anonymous):

how do you convert 3r=cos(theta) to rectangular form??

OpenStudy (jdoe0001):

do you know what a perfect square trinomial is?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

well... then you may want to cover your perfect trinomials sections first,... since to make it rectangular, you'd need it -> http://www.youtube.com/watch?v=xGOQYTo9AKY

OpenStudy (anonymous):

ok, I've been working on it for about 10 mins now, does the answer come out to 3x^2+3y^2-x=0????

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

\(\bf 3r=cos(\theta)\quad \textit{multiply both sides by "r"}\\ \quad \\ 3r^2=rcos(\theta)\implies 3(x^2+y^2)=x\implies 3x^2+3y^2-x=0\)

OpenStudy (anonymous):

thanks a lot man

OpenStudy (jdoe0001):

then you'd need to "completing the square" to get the actual equation, which is just a circle